# Opponent Shaping for Antibody Development

### v6

@@ 2.1.1 Virus-Antibody Game
To meet the computational demands of our opponent shaping approach, which requires rapid evaluation of numerous antibody-antigen interactions, we reimplement the core binding calculation of Absolut! using JAX [13], a framework that facilitates GPU-accelerated computation, see Figure 1c. Our efficient JAX implementation and the GPU acceleration results in a 10,000-fold speedup compared to the original implementation, see Table 1.
@@
Table 1: Speed comparison of the binding calculation for a single antibody-antigen query between the original implementation of Absolut! [11] and our reimplementation of Absolut! in JAX [13].
2.1.2 Simulated Viral Escape via Evolution
Viral Escape is the process of a virus adapting to evade the host’s immune system [12]. From a game theory perspective, this is the process of the virus “learning to play the game”. To simulate the space of possible viral escape trajectories against a chosen antibody, we employ an evolutionary algorithm. This approach approximates the process of viral evolution, where mutations that increase fitness (in our case, the virus’s payoff $R_{v}$) are more likely to persist, propagate, and become an ancestor of future generations.
We simulate this process of viral escape for a specified horizon of $H$ steps, with each step representing a single generation of viral evolution. This is the target duration of therapy efficacy. For each independent viral trajectory, we use the initial virus strain $v$ to generate a population of $P$ virus variants through random mutations. These are point mutations in the amino acid sequence of the viral antigen and they are introduced according to a predefined mutation rate. Each variant’s fitness is evaluated based on its payoff $R_{v}$ against the fixed antibody $a$. Importantly, this process captures the myopic nature of viral evolution: the virus optimises its fitness only with respect to the current antibody, unable to anticipate future human interventions or antibody changes. The next generation of viruses is then created by probabilistically selecting the variants based on their fitness, with fitter variants more likely to be chosen for replication and subsequent mutation. This process repeats for $H$ generations, creating one trajectory of viral escape, see the Simulated Viral Escape via Evolution in Figure 1a and Methods Section 4.2 for more details.
@@ 2.1.2 Simulated Viral Escape via Evolution
As an example implementation of our method we use the antigen protein from the Dengue Virus for all our experiments, specifically, the structure with Protein Data Bank (PDB) code 2R29 [21, 22]. We discretise this structure using the Absolut! framework to generate the docking poses, which we then use to calculate the binding strength $B$ of any antibody-virus pair using our JAX implementation. In the viral escape step, we mutate only the amino acid sequence of the dengue envelope antigen, which is composed of 97 amino acids, and do not consider other components of Dengue Virus. Importantly, while we alter the amino acid sequence, we assume the structure of the antigen does not significantly change for the calculation of binding strength.
2.1.3 Optimisation of Antibody Shapers
Our antibody optimisation process aims to generate antibodies that can prevent viral escape. In particular, we want antibodies which maintain their performance even once the virus has had the opportunity to evolve in response to them. For this purpose, we define the antibody fitness function $F^{H}_{v}(a)$ as the antibody’s average performance over multiple simulated viral escape trajectories of virus $v$:
$F^{H}_{v}(a)=\mathop{\mathbb{E}}_{\eta\text{ escape trajectories of }v}\left[\mathop{\mathbb{E}}_{H\text{steps of escape}}(R_{a}(v^{\prime},a))\right]$ (2)
Specifically, for each of the $\eta$ trajectories, we simulate the viral escape process for $H$ steps and calculate the average antibody payoff $R_{a}$ over these steps, where $v^{\prime}$ represents the possible mutated variants of the original virus $v$. The final fitness is the average over these $\eta$ simulations, i.e. $\eta$ Monte Carlo roll-outs. The fitness function $F^{H}_{v}(a)$ approximates the antibody’s long-term effectiveness against evolving viral populations.
We then employ an evolutionary algorithm on the antibodies themselves to optimise this fitness function. In particular, starting from random antibodies at step 00, optimisation proceeds iteratively for $N$ steps where we select the antibody with the highest fitness at each step and generate a new population of $P_{a}$ antibodies by introducing single mutations to the selected antibody. This process of evaluation, selection, and mutation continues, gradually improving the antibody’s robustness against, and ability to limit, viral escape. See Antibody Optimisation Loop in Figure 1a and Methods Section 4.3 for more details.
We refer to the resulting optimised antibodies as shapers, or more specifically as $H=5$ shapers or $H=100$ shapers, reflecting the horizon length of the viral escape simulations used during their optimisation. Myopic antibodies are antibodies where horizon length $H=0$ and thus $F^{0}_{v}(a)=R_{a}(v,a)$. As we will discuss in more detail, the choice of $H$ significantly impacts the antibody’s long-term effectiveness at the cost of computational resources required for optimisation [^1] A single step of antibody optimisation requires simulating $P_{a}\times\eta$ independent viral escapes, each for horizon $H$ steps. However, we can parallelise over both $P_{a}$ and $\eta$., making horizon $H$ a vital parameter to consider.
From a machine learning perspective, this process can be viewed as meta-optimisation or meta-learning. The inner loop of our optimisation is the viral escape simulation, where we model the virus learning to evade the current antibody. The outer loop or meta-loop is the antibody optimisation process itself, which learns to generate antibodies that perform well across many possible viral escape trajectories. We are designing antibodies that anticipate and influence the escape process of the virus, embodying the core principle of opponent shaping in this biological context.
2.2 Antibody Shapers Minimise Viral Escape
2.2.1 Shapers vs. Myopic Antibodies
@@ 2.2 Antibody Shapers Minimise Viral Escape
Figure 2: Optimised antibody shapers outperform myopic antibodies. a Distribution of antibody shapers (in orange) which were optimised with horizon $H=100$ vs. myopic antibodies distribution (in blue). We highlight the top 10% shapers with respect to $F^{100}_{v}(a)$ in red and the top 10% myopic antibodies with respect to $R_{a}(v,a)$ in green. The x-axis is the myopic antibody fitness $R_{a}(v,a)$ and the y-axis is the escape averaged antibody fitness over 100 steps of viral escape. Higher values on both axes indicate better performance. b Viral escape curves for different steps of the antibody optimisation process for antibody shapers optimised with horizon 100 (solid lines) and myopic antibodies (dashed lines). The lighter lines indicate early antibody optimisation steps and the darker lines show the later steps. The x-axis shows the evolutionary steps of viral escape. The y-axis represents the virus fitness/payoff $R_{v}(v,a)$, where higher values indicate better virus fitness (and lower values denote better antibody performance). c, d Antibody optimisation learning curves for a varying horizon length. The x-axis shows the antibody optimisation steps (c) or the number of samples from the binding simulator (d). The y-axis represents the escape averaged antibody fitness over 100 steps of viral escape $F^{100}_{v}(a)$. Error bars correspond to the standard error. Higher values indicate better performance.
Our results demonstrate a clear advantage of shapers in the escape-averaged objective $F^{100}_{v}(a)$. The mean of the shapers distribution significantly exceeds that of the myopic distribution, as evident from the marginal density plot in Figure 2a. Notably, none of the myopic antibodies outperforms the top 10% of shapers in this long-term objective. However, there is a trade-off between short-term and long-term optimisation. While shapers do better on the escape-averaged objective, they underperform compared to myopic antibodies in optimising the immediate payoff $R_{a}(v,a)$ against the initial, non-mutated virus $v$.
@@
At the outset of the antibody optimisation process (step 00) both the shapers and the myopic antibodies induce similar escape curves, an expected outcome given their initialisation from random antibody sequences. However, as we examine antibodies from later optimisation steps, we observe diverging trends. Myopic antibodies cause the viral fitness to be lower in the initial escape steps, outperforming the shapers. After about $10$ escape steps, corresponding to $\approx 10$ viral mutations, the two antibody types perform similarly. Beyond that, shapers demonstrate superior results in later escape stages, more effectively preventing viral escape.
@@
2.2.2 Antibody Shapers with Varying Horizons
@@
To illustrate this trade-off, we conduct an additional experiment shown in Figure 2d. Here, instead of fixing the number of optimisation steps $N$, we constrain the total number of binding samples - queries to our binding strength simulator used to evaluate all antibody and virus payoffs throughout the optimisation process - to be constant across different horizons. This approach provides a performance comparison that accounts for the computational resources necessary across varying horizon lengths. Interestingly, $H=20$ shapers perform strongly, nearly matching the performance of those optimised with horizon $H=100$ for a given number of antibody optimisation steps, and far exceeding it when accounting for the differing computational cost. This suggests that using a cheaper, shorter-horizon, proxy for the true antibody objective $F^{100}_{v}(a)$ can yield substantial benefits.
More generally, we find that the optimisation horizon significantly influences the performance of antibody shapers. While longer horizons generally lead to better long-term performance, the optimal horizon length is dependent on the available computational resources. Thus, it is important to consider the balance between computational cost and the fidelity of the optimisation objective when designing antibodies for long-term effectiveness against evolving viral populations.
2.3 How the Antibody Shapers Prevent Viral Escape
2.3.1 Attack is the Best Defence
@@
Figure 3: Effect of myopic and shaper antibodies on viruses induced by other antibodies. We optimise $80$ different antibodies $a_{H}$ for each horizon (Myopic, $H=5,H=10,H=20,H=100$), these are represented by the y-axis. We simulate the viral escape to each of these antibodies for $100$ steps and we group the escape viruses $v_{H}$ by the horizon $H$ of the antibody that induced them, these escape viruses are represented by the x-axis. In colour, we show the mean antibody payoff $R_{a}(v_{H},a_{H^{\prime}})$ for each group of optimised antibodies $a_{H^{\prime}}$ against the final escape viral variant $v_{H}$ induced by other antibodies optimised with horizon $H$. Darker colours correspond to better antibody payoff.
@@
To disentangle these strategies, we seperately evaluate the antibodies and the viruses that evolve in response to them. To do this, we compare the viruses against other antibodies which did not influence the viral evolution. The intuition is that an antibody which is good at shaping (good attack), but less robust (poor defence), will induce viruses which other antibodies will perform well against. Specifically, we generate antibodies $a_{H}$ for each horizon $H$ and simulate viral escape against these antibodies for 100 steps, resulting in viruses $v_{H}$. For all pairs of horizons $(H,H^{\prime})$, we then cross-evaluate the antibody payoff $R_{a}(v_{H},a_{H^{\prime}})$. Figure 3 presents the result of this analysis.
@@
Interestingly, viruses $v_{100}$ induced by $H=100$ shapers are consistently more exploitable by antibodies across all optimisation horizons. This suggests that $H=100$ shapers actively shape the escape trajectories of the virus in a way that makes the resulting variants more susceptible to antibody binding in general. However, this shaping effect comes at a cost. The $H=100$ shapers ($a_{100}$) show slightly lower payoffs compared to the peak performance of shorter-horizon antibodies ($a_{5}$ and $a_{10}$) against the viruses $v_{100}$ induced by the $H=100$ shapers (see rightmost column of Figure 3). This trade-off indicates that to exert a stronger shaping influence on viral evolution, $H=100$ shapers sacrifice some degree of immediate performance or robustness, that is their ability to perform well against a wide range of viruses. Therefore, a potential strategy could involve using a “mixture” of antibodies as therapy: some optimised for shaping the virus’s evolutionary trajectory, and others designed for strong immediate binding.
2.3.2 Amino Acid Distribution in Shapers and Myopic Antibodies

### v8

@@ 5.1 Absolut! Speedup
To meet the computational demands of our opponent shaping approach, which requires rapid evaluation of numerous antibody-virus interactions, we reimplement the core binding calculation of Absolut! (Robert et al., 2022) using JAX (Bradbury et al., 2018), a framework that facilitates GPU-accelerated computation (Figure 2c). Our efficient JAX implementation and the GPU acceleration results in a 10,000-fold speedup compared to the original implementation, see Table 1.
@@
Table 1: Comparison of the time taken to compute a single binding query between the original implementation of Absolut! (Robert et al., 2022) and our reimplementation of Absolut! in JAX (Bradbury et al., 2018). The original Absolut! implementation runs on CPU only, hence the difference in evaluation hardware.
5.2 Dengue Virus
We use the antigen protein from the Dengue Virus for our main experiments, specifically, the structure with Protein Data Bank (PDB) code 2R29 (Berman et al., 2000; Lok et al., 2008). First, Absolut! processes this structure to generate binding-relevant information, which is then used by our JAX implementation (details given in Appendix D).
Figure 3: Shapers Outperform Myopic Antibodies. a Distribution of antibody shapers optimised with horizon $H=100$ (orange) vs. myopic antibodies distribution (blue). We highlight the top 10% shapers with respect to $F^{100}_{v}(a)$ in red, and the top 10% myopic antibodies with respect to $R_{a}(v,a)$ in green. The x-axis is the myopic antibody fitness $R_{a}(v,a)$ and the y-axis is the escape averaged antibody fitness for $H=100$, i.e., $F_{\hat{v}}^{100}(a)$. Higher values on both axes indicate better performance. b Viral escape curves (inner loop performance) for different steps of the antibody optimisation process (outer loop) for antibody shapers optimised with horizon 100 (solid lines) and myopic antibodies (dashed lines). The lighter lines indicate early antibody optimisation steps, and the darker lines show the later steps. The x-axis shows the evolutionary steps of viral escape. The y-axis represents the virus fitness/payoff $R_{v}(v,a)$, where higher values indicate better virus fitness (and lower values denote better antibody performance, so lower is better for us). c, d Antibody optimisation learning curves (outer loop performance) for a varying horizon length. The x-axis shows the antibody optimisation steps, i.e., meta-steps (c) or the number of samples from the binding simulator (d). The y-axis shows antibody fitness $F^{100}_{v}(a)$. Error bars correspond to the standard error. Higher values indicate better performance.
@@ 5.2 Dengue Virus
In the viral escape step, we mutate only the amino acid sequence of the dengue envelope antigen, which is composed of $N_{v}=97$ amino acids, and do not consider other components of Dengue Virus. Importantly, we assume that the structure of the antigen does not significantly change over the course of viral escape. All experiments but the ones we discuss in Section 6.3 and Appendix A use the dengue virus.
5.3 Additional Viruses and Bacterium
To demonstrate robustness of the experimental results we achieve with ADIOS on dengue virus we conduct additional experiments with three other viruses and one bacterium. The three viral antigens we use are: West Nile Virus, PDB code 1ZTX (Nybakken et al., 2005); Influenza Neuraminidase Virus, PDB code 4QNP (Wan et al., 2015) and MERS-CoV Virus with PDB code 5DO2 (Li et al., 2015). Furthermore, we show that ADIOS can be easily applied to other pathogens, such as bacteria, too. We perform an extra experiment with the Clostridium Difficile Bacterium, PDB code 4NP4 (Orth et al., 2014).
6 Results
6.1 Shapers vs. Myopic Antibodies
Figure 4: Shapers Outperform Myopic Antibodies on Other Viruses and a Bacterium. Antibody optimisation learning curves (outer loop performance) for varying horizon lengths across three viruses - West Nile Virus, Influenza Virus, and MERS-CoV - as well as the bacterium Clostridium Difficile. The x-axis shows the antibody optimisation steps, i.e., meta-steps and the y-axis shows the antibody fitness $F^{100}_{v}(a)$. Raw fitness values are dependent on Absolut! scores and are relative to the specific antigen, i.e., absolute values should not be compared between different viruses or bacterium but rather the overall trends. Error bars correspond to the standard error. Higher values indicate better performance. Full set of ADIOS results for these four pathogens is provided in Appendix A.
@@ 6 Results
Our results demonstrate a clear advantage of shapers in the escape-averaged objective $F^{100}_{v}(a)$. The mean of the shapers distribution significantly exceeds that of the myopic distribution, as evident from the marginal density plot in Figure 3a. Notably, none of the myopic antibodies outperform any of the top 10% of shapers in this long-term objective. However, there is a trade-off between short-term and long-term optimisation. While shapers do better on the escape-averaged objective, they underperform on the myopic objective $R_{a}(v,a)$.
@@
At the outset of the antibody optimisation process (step 00), both the shapers and the myopic antibodies induce similar escape curves, an expected outcome given their initialisation from random antibody sequences. However, as we examine antibodies from later optimisation steps, we observe diverging trends. Myopic antibodies cause the viral fitness to be lower in the initial escape steps, outperforming the shapers. After about $10$ escape steps, corresponding to $\approx 10$ viral mutations, the two antibody types perform similarly. Beyond that, shapers demonstrate superior results in later escape stages, more effectively preventing viral escape.
@@
6.2 Antibody Shapers with Varying Horizons
@@
To illustrate this trade-off, we conduct an additional experiment shown in Figure 3d. Here, instead of fixing the number of optimisation steps $N$, we constrain the total number of binding samples - queries to our binding strength simulator used to evaluate all antibody and virus payoffs throughout the optimisation process - to be constant across different horizons. This approach provides a performance comparison that accounts for the computational resources necessary across varying horizon lengths. Interestingly, $H=20$ shapers perform strongly, nearly matching the performance of those optimised with horizon $H=100$ for a given number of antibody optimisation steps, and far exceeding it when accounting for the differing computational cost. This suggests that using a cheaper, shorter-horizon proxy for the true antibody objective $F^{100}_{v}(a)$ can yield substantial benefits.
More generally, we find that the optimisation horizon significantly influences the performance of antibody shapers. While longer horizons lead to better long-term performance, the optimal horizon length is dependent on the available computational resources. Thus, it is important to consider the balance between computational cost and the fidelity of the optimisation objective when designing antibodies for long-term effectiveness against evolving viral populations.
6.3 Antibody Shapers for Other Viruses and Bacterium
To test whether the shaping effects observed on dengue generalise, we evaluate ADIOS on three additional viruses: West Nile, Influenza, and MERS-CoV; as well as the Clostridium Difficile bacterium (Figure 4, Appendix A). In all cases, we observe consistent trends: shapers outperform myopic antibodies in escape-averaged fitness, confirming that ADIOS can successfully achieve shaping across diverse pathogens.
For West Nile virus and MERS-CoV, $H=100$ shapers appear to perform worse than shorter-horizon $H=20$ antibodies, see Figure 4. However, given that all antibodies are only optimised for 30 meta-steps, and the compute-normalised (bottom row) plots in Figure A.1 show clearly that $H=100$ shapers have not yet converged, we hypothesise that $H=100$ shapers would ultimately yield the best performance if given more optimisation time.
Interestingly, the shaping effect is especially strong on the Clostridium difficile bacterium, where $H=100$ shapers significantly outperform all other antibody types across all reported metrics; see right-most row in Figure A.1. These results suggests that different antigens can exhibit very different behaviour in a shaping setting. However, ADIOS is able to produce antibodies with effective shaping behaviour across them.
6.4 Attack is the Best Defence
@@
Figure 5: Robustness vs. Shaping. We optimise $80$ different antibodies $a_{H}$ across multiple horizons (Myopic, $H=5,H=10,H=20,H=100$), these are represented by the y-axis. We simulate the viral escape to each of these antibodies for $100$ steps, and we group the escape viruses $v_{H}$ by the horizon $H$ of the antibody that induced them; these escape viruses are represented by the x-axis. In colour, we show the mean antibody payoff $R_{a}(v_{H},a_{H^{\prime}})$ for each group of optimised antibodies $a_{H^{\prime}}$ against the final escape viral variant $v_{H}$ induced by other antibodies optimised with horizon $H$. Darker colours correspond to better antibody payoff.
@@
To disentangle these strategies, we separately evaluate the antibodies and the viruses that evolve in response to them. To do this, we compare the viruses against other antibodies which did not influence the viral evolution. The intuition is that an antibody which is good at shaping (good attack), but less robust (poor defence), will induce viruses which other antibodies will perform well against. Specifically, we generate antibodies $a_{H}$ for each horizon $H$ and simulate viral escape against these antibodies for 100 steps, resulting in viruses $v_{H}$. For all pairs of horizons $(H,H^{\prime})$, we then cross-evaluate the antibody payoff $R_{a}(v_{H},a_{H^{\prime}})$. Figure 5 presents the result of this analysis.
@@
Interestingly, viruses $v_{100}$ induced by $H=100$ shapers are consistently more exploitable by antibodies across all optimisation horizons. This suggests that $H=100$ shapers actively shape the escape trajectories of the virus in a way that makes the resulting variants more susceptible to antibody binding in general. However, this shaping effect comes at a cost. The $H=100$ shapers ($a_{100}$) show slightly lower payoffs compared to the peak performance of shorter-horizon antibodies ($a_{5}$ and $a_{10}$) against the viruses $v_{100}$ induced by the $H=100$ shapers (see rightmost column of Figure 5). This trade-off indicates that to exert a stronger shaping influence on viral evolution, $H=100$ shapers sacrifice some degree of immediate performance or robustness, that is, their ability to perform well against a wide range of viruses. Therefore, a potential strategy could involve using a mixture of antibodies as therapy, where some are optimised for shaping the virus’s evolutionary trajectory, and others are designed for strong immediate binding.
Figure 6: Shaping with External Pressure. A similar experiment to Figure 5, but with the additional external pressure of a separate myopic antibody (see Equation 3).
To investigate whether shaping persists in more realistic scenarios with multiple therapeutic pressures, we conduct an additional experiment. Using the same groups of antibodies $a_{H}$, we simulate viral escape with external pressure from a myopic antibody $a_{Ext}$ (not included in the original $a_{myopic}$ set). The viruses $v_{H+Ext}$ now evolve according to a modified payoff:
$R_{v}^{Ext}(v,a,a_{Ext})=\frac{1}{2}R_{v}(v,a)+\frac{1}{2}R_{v}(v,a_{Ext})$ (3)
which represents the scenario where multiple therapies are present in the environment (for example, during COVID-19 when multiple vaccines were available). Figure 6 shows that while the shaping effect is somewhat reduced compared to our original results, it remains clearly visible. This demonstrates that our shaping approach transfers to test regimes where external pressures from other therapies are present.
6.5 Explainability Analysis
To understand what distinguishes shapers from myopic antibodies, we conduct two complementary analyses of amino acid distributions and binding poses (Appendices B and C). Examining amino acid distributions, we find that long-horizon shapers exhibit more uniform distributions, while myopic antibodies tend to cluster around amino acids with extreme binding energies. We hypothesise that by maintaining diversity in their amino acid composition, shapers can preserve robustness against viral mutation since the virus cannot easily escape by avoiding specific, high-binding, parts of the antibody.
Through analysis of binding using pose matrices, representing which parts of the antibody and virus bind with each other, we observe these interaction patterns significantly change as the virus adapts. However, the type of antibody influences the nature of these changes: $H=100$ shapers actively constrain viral evolution by both preventing unfavourable binding configurations and preserving favourable ones. While these findings are specific to our Absolut! binding simulator, they hint at explainable strategies that shapers use to influence viral evolution, which could inform future antibody design approaches.
7 Conclusion & Future Work
In this work, we introduce ADIOS, a meta-learning framework for designing therapeutic antibodies that not only defend against current viral strains but instead actively shape viral evolution. We provide a GPU-accelerated JAX implementation of Absolut!, enabling rapid simulation of viral escape trajectories and outer-loop optimisation. Our results demonstrate that shapers are not only more robust against viral escape, but they also shape viral evolution toward more targetable variants. Lastly, we provide an explainability analysis of how shapers achieve this level of robustness and influence, which we hope will inspire practitioners.
Although dengue virus served as our primary benchmark, we evaluated ADIOS on three other viruses (West Nile, Influenza and MERS-CoV) and on the bacterium Clostridium Difficile. In all four cases ADIOS consistently achieves shaping. These results confirm that ADIOS can generalise across a diverse set of pathogens. More broadly, the same opponent-shaping principle can be transferred to monoclonal-antibody (mAb) therapy for cancer (Zahavi & Weiner, 2020). In that setting, the outer loop would optimise therapeutic mAbs, while the inner loop would simulate the evolution of cancer-cell growth-factor receptors; the goal would be to shape cancer cells into cells that do not proliferate well. Exploring this direction, together with other bacterial or antimicrobial-resistance scenarios, remains an exciting avenue for future work.
While our current implementation uses simplified binding and evolutionary escape models that prevent direct therapeutic application, ADIOS could be integrated with more sophisticated models, like AlphaFold3 (Abramson et al., 2024), to better capture evolving viral and antibody structures. As computational models of protein interactions and evolutionary processes continue to improve, ADIOS has the potential to transform how we develop therapies against viruses, cancers, and other evolving adversaries.
Acknowledgements
We would like to thank Marius Urbonas, Duygu Açıkalın, Michael Matthews, Benjamin Ellis, Benedetta L. Mussati, Mattie Fellows, Lisa Zillig, Ting Lee, Matthew Raybould and Charlotte Deane for their invaluable contributions to this work. Their thoughtful insights helped guide the project direction, their detailed feedback on earlier drafts significantly enhanced the manuscript’s clarity and accessibility, and their support during the review process was instrumental in addressing reviewers’ concerns. We also thank the anonymous reviewers for their constructive feedback, which led to improvements in the paper. This work was supported by UK Research and Innovation and the European Research Council - Jakob Foerster is partially funded by the UKRI grant EP/Y028481/1, originally selected for funding by the ERC. This work was also supported by Exscientia, which provided Aleksandra Kalisz with a PhD studentship.
Impact Statement
This work introduces a machine learning framework for antibody design that could contribute to developing more effective therapies against evolving pathogens like viruses. While this has a potential positive societal impact through improved disease treatment and pandemic preparedness, we acknowledge that therapeutic development tools carry risks if misused. Our current implementation uses simplified models, and significant additional research and safety validation would be required before any real-world therapeutic applications.
Appendix A Experimental Results on Other Viruses and Bacterium
Figure A.1: Shapers Outperform Myopic Antibodies on Other Viruses and a Bacterium. The figure shows results across four different pathogens: West Nile Virus, Influenza Virus, MERS-CoV Virus, and Clostridium Difficile Bacterium (left to right columns). First row Distribution of antibody shapers optimised with horizon $H=100$ (orange) vs. myopic antibodies distribution (blue). We highlight the top 10% shapers with respect to $F^{100}_{v}(a)$ in red, and the top 10% myopic antibodies with respect to $R_{a}(v,a)$ in green. The x-axis is the myopic antibody fitness $R_{a}(v,a)$ and the y-axis is the escape averaged antibody fitness for $H=100$, i.e., $F^{100}_{v}(a)$. Higher values on both axes indicate better performance. Second row Viral escape curves (inner loop performance) for different steps of the antibody optimisation process (outer loop) for antibody shapers optimised with horizon $H=100$ (solid lines) and myopic antibodies (dashed lines). The lighter lines indicate early antibody optimisation steps, and the darker lines show the later steps. The x-axis shows the evolutionary steps of viral escape. The y-axis represents the virus fitness/payoff $R_{v}(v,a)$, where higher values indicate better virus fitness (and lower values denote better antibody performance, so lower is better for us). Third row Antibody optimisation learning curves (outer loop performance) for varying horizon lengths. The x-axis shows the antibody optimisation steps, i.e., meta-steps, and the y-axis shows antibody fitness $F^{100}_{v}(a)$. Forth row Antibody optimisation learning curves accounting for computational cost. The x-axis shows the number of samples from the binding simulator, and the y-axis shows antibody fitness $F^{100}_{v}(a)$. Higher values indicate better performance. Error bars correspond to the standard error. In all these results, raw payoff/fitness values are dependent on Absolut! scores and are relative to the specific antigen, i.e., absolute values should not be compared between different viruses or bacterium but rather the overall trends.
Appendix B Amino Acid Distribution in Shapers and Myopic Antibodies